T2(N) = Θ(N) Which algorithm is the least preferred one? a) Algorithm 1 b) Algorithm 2 ) Algorithm 3 Algorithm 1 and 3 2. Given the following function: (10 points) int F(int N) if (N0lIN1) return 2; else return N *FON-2); Show steps to find out F(5) Also give the runtime
Answer #1
F(5)
= 5 * F(3)
= 5 * 3 * F(1)
= 5 * 3 * 2
= 5 * 6
= 30

so, F(5) is 30

runtime complexity is O(n)

Please let me know i f you need me to make anu changes